OpenStudy (anonymous):

Solve the formula for the given variable. Assume all variables represent nonnegative numbers. A=14s^2, for s

OpenStudy (anonymous):

A/12 = s^2 \[\sqrt{A/12} = s\]

OpenStudy (anonymous):

thank yu

OpenStudy (anonymous):

thanks for the medal it can actually simplfy a bit more to \[\sqrt{A}/2\sqrt{3}\] or \[\sqrt{3}\sqrt{A}/6\] = \[\sqrt{3A}/6\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how do you construct natural numbers from successor?